what is 3(2n-3)+6n-2<28
1
Expert's answer
2019-10-11T10:08:58-0400

"3(2n-3)+6n-2<28\\
6n-9 + 6n-2<28\\
12n-11<28\\
12n<28+11\\
12n<39\\
4n<13\\
n<\frac{13}{4} =3\frac{1}{4}\\
Answer: n<3\frac{1}{4}"